OpenStudy (samigupta8):

the interval for which the series 1+(x-1)+(x-1)^2+.......∞ may be summed is

OpenStudy (samigupta8):

plss....tell without the use of limit

OpenStudy (raden):

that series would be convergen if -1 < r < 1

OpenStudy (raden):

with r = a2/a1 = (x-1)/1 = x-1 so, -1 < x-1 < 1 add by 1 , we get -1 + 1 < x-1+1 < 1+1 0 < x < 2

OpenStudy (raden):

that interval makes the series can be summed

OpenStudy (samigupta8):

if the value of r lies between -1 and 1 then only it is possible to sum it up is this a fix case for it

OpenStudy (raden):

yes, that's the rule if given geometri series and so that it converges the ratio must be in that interval